OpenStudy (kanwal32):

if for positive integer r>1,n>2 the coefficients of the (3r)th and (r+2)th power of x in the expression of (1+x)^2n are equal then A)2r=n B) n=3r C)n=2r+1 D)n=2r-1

OpenStudy (kanwal32):

i got it A pls anyone can check @ShadowLegendX @Directrix @ganeshie8 @Kainui

OpenStudy (vishweshshrimali5):

Can you expand the expression and write the general term?

OpenStudy (kanwal32):

2nC3r-1=2nCr+1

OpenStudy (kanwal32):

3r-1=2n-r-1

OpenStudy (vishweshshrimali5):

Or, 3r - 1 = r + 1 is also possible... but that would mean r = 1 which is not true...

OpenStudy (vishweshshrimali5):

Great after that ?

OpenStudy (kanwal32):

n=2r

OpenStudy (kanwal32):

is it right

OpenStudy (vishweshshrimali5):

Yeps! Great job!